### v5

@@ 3 Rank-minimizing Mechanism
That is, if a rank-minimizing mechanism is employed and $a_{1}$ chooses either $\alpha$ or $\alpha^{\prime}$, then $a_{1}$ must have either $o_{1}$ or $o_{2}$. Thus, the only strategic interest of $a_{1}$ is increasing the probability to obtain $o_{1}$. If $f^{\ast}$ is employed, then $\alpha^{\prime}=r_{a_{1}}$ is strategically dominated by $\alpha$ for $a_{1}$ due to (2). Therefore, we conclude that even if a rank-minimizing mechanism that satisfies weak ETE is employed, truth-telling may be strategically dominated by some other strategy. Note that in the same case, $f^{\ast}$ does not satisfy ETE. First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing$,∅ , as long as the assignment is not wasteful. Thus, $\alpha$ and $\alpha^{\prime}$ are essentially the same in this case. However, $\left(f^{\ast}\left(\alpha^{\prime},\alpha,\alpha\right)\right)_{a_{1}o_{1}}=0,$ but $\left(f^{\ast}\left(\alpha^{\prime},\alpha,\alpha\right)\right)_{a_{1}o_{2}}>0$ or $\left(f^{\ast}\left(\alpha^{\prime},\alpha,\alpha\right)\right)_{a_{1}o_{3}}>0$ must be satisfied. Hence $f^{\ast}$ does not satisfy ETE.
@@ 4 Refusal option
In this section, we consider the following three-stage game, which is a modification of that in Section 3; that is, the first and second stages in this game are the same as in the previous game. In the third stage, all agents decide whether to accept or refuse the assigned object type. If an agent refuses their assignment, then the agent instead obtains the null object type $\varnothing$ (the outside option). That is, in this game, the null object type has two characteristics: first, there are sufficiently many copies of this type to allocate to all individuals, and second, agents can always obtain it at no cost as long as they abandon their assignment.
This third stage is also considered by Featherstone (2020) but is not explicitly addressed by almost other previous studies related to ours. [^4] As exceptions, Feigenbaum et al. (2020) and Afacan (2022) consider an assignment problem where, after the initial allocation is decided, if someone refuses the assigned object, it is reallocated. Moreover, Doğan and Yenmez (2019) also explicitly consider this stage. However, this stage is implicitly considered by many of them; that is, they consider an ex-post individually rational (or more strictly an ex-post efficient) matching as a requirement and if an assignment is ex-post individually rational, no agent refuses their assignment. See, for example, Bogomolnaia and Moulin (2001) on the property.
@@
$g\left(x,r\right)_{ao}$ $=$ 0 for all $0\text{ for all }R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$
@@
We reconsider strategic problems of a fair rank-minimizing mechanism in this three-stage game. For a mechanism $f$ and an agent $a$, (revealing) $p_{a}^{\prime}$ is said to be weakly strategically dominated by (revealing) $p_{a}$ with the refusal option if $a$ weakly prefers $g\left(f\left(p_{a},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$. Moreover, for $f$ and $a$, (revealing) $p_{a}^{\prime}$ is said to be (strictly) strategically dominated by (revealing) $p_{a}$ with the refusal option if $p_{a}^{\prime}$ is weakly strategically dominated by $p_{a}$ with the refusal option and there exists some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $g\left(f\left(p_{a},p_{-a}^{\prime}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}^{\prime}\right),r\right)$.
@@
We show that, under the uniform rank-minimizing mechanism, truth-telling is strategically dominated by another strategy when the agents have the refusal option. We define the following specific strategies. Let $\mathcal{D}_{a}\subseteq\mathcal{P}$ be a set of preferences such that $d_{a}\in\mathcal{D}_{a}$ if (i) $R\left(o,r_{a}\right)=R\left(o,d_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and (ii) $R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$. First, the ranks of the truly acceptable object types are unchanged. Second, the null object type $\varnothing$ is the least preferable one under $d_{a}\in\mathcal{D}_{a}$. We call the strategy of agent $a$ of revealing a preference in $\mathcal{D}_{a}$ an outside option demotion strategy (hereafter ODS) of agent $a$. Under an ODS, every object type that is ranked below $\varnothing$ in the agent’s true preference is instead ranked above $\varnothing$ in $d_{a}$. Note that an ODS (revealing $d_{a}\in\mathcal{D}_{a}$) increases the number of acceptable object types revealed by $a$; that is, $R\left(\varnothing,r_{a}\right)\leq R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$ is satisfied.
@@
Featherstone (2020) defines the full extension of $r_{a}$denoted by $e_{a}$ in this study such that (i) $R\left(o,r_{a}\right)=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (ii) $R\left(\varnothing,e_{a}\right)=\left|\mathcal{O}\right|$, and $R\left(o,r_{a}\right)-1=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$; that is, $e_{a}\in\mathcal{D}_{a}$ is such that all non-null object types are placed in their true preference orders. Revealing the full extension of $r_{a}$ is one of the ODSs of $a$.
@@
Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their first-best object $o_{1}$. First, under $\left(r_{a_{1}},\beta,\beta^{\prime}\right),$varnothing is ranked lowest by $a_{2}$ and $a_{3}$ but $a_{1}$ ranks $\varnothing$ as their second-best. Thus, in any rank-minimizing mechanism, $a_{1}$ is assigned $\varnothing$ with probability 1. On the other hand, since $a_{1}$ places $\varnothing$ at the bottom of their ranking. As a result. under $\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$, $a_{1}$ is assigned $o_{1}$—the object that is truly their best—with probability $1/2$, and $\varnothing$ with probability $1/2$. Thus, in this case, $a_{1}$ is better off by choosing $d_{a_{1}}$ instead of $r_{a_{1}}$.
@@
$f^{U}\left(r_{a_{1}},\beta,\beta\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1/2&1/2&0\\ 1/2&1/2&0\end{array}\right),$
$f^{U}\left(d_{a_{1}},\beta,\beta\right)=\left(\begin{array}[]{ccc}1/3&1/3&1/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$
@@
$g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$
@@
and $\text{and }g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)_{a^{\prime}\varnothing}$ $=$ $\frac{1}{3}>0,$
@@
If $a_{1}$ is assigned to $o_{2}$, then either $a_{2}$ or $a_{3}$ will be assigned to $\varnothing$. However, in this case, $a_{1}$ will refuse $o_{2}$ and thus $o_{2}$ will be remained unallocated. In this case, rather than leaving $o_{2}$ unallocated, it is more efficient to assign $o_{2}$ to the agent ($a_{2}$ or $a_{3}$) who would otherwise receive $\varnothing$ and instead assign $a_{1}$ to $\varnothing$ from the beginning. This adjustment ensures that $o_{2}$ is allocated efficiently. Hence revealing $d_{a_{1}}$ and refusing an unacceptable assignment may benefit $a_{1}$ but lead to a wasteful allocation, as some agents may prefer the object type that $a_{1}$ refuses.
Note that $\left(d_{a_{1}}\left(=\beta\right),\beta,\beta\right)$ can be a Nash equilibrium. Trivially, $a_{1}$ has no incentive to change $d_{a_{1}}$. On the other hand, suppose that both $a_{2}$ and $a_{3}$ prefer the assignment such that each object $o\in\left\{o_{1},o_{2},\varnothing\right\}$ assigns with the probability $1/3$ to the assignment such that $o_{2}$ assigns with probability $1$. This ensures that neither $a_{2}$ nor $a_{3}$ has an incentive to change $\beta$ and $\beta$ to $(o_{2},o_{1},\varnothing)$ or $(o_{2},\varnothing,o_{1})$. Then, in this case, $\left(d_{a_{1}},\beta,\beta\right)$ is a Nash equilibrium. Thus, a wasteful assignment can occur as a Nash equilibrium of agents’ strategies when the assignment is determined by the uniform rank-minimizing mechanism.
@@
Suppose that the uniform rank-minimizing mechanism is employed. Then, for all agents, truth-telling is weakly strategically dominated by any ODS with the refusal option.
@@
Theorems 1 and 2 do not imply that an ODS is a dominant strategy. However, if there is $a\in\mathcal{A}$ whose true preference satisfies (3) for some $o,o^{\prime}\in\mathcal{O}$, since truth-telling is strategically dominated by another strategy, $a$ has a strong incentive to manipulate its preference ranking, while one of the ODSs may not necessarily be chosen.
As stated earlier, multiple ODSs may exist. Even if an ODS strategically dominates truth-telling, another ODS that does not strategically dominate truth-telling may also exist. In such a case, the latter ODS is strategically dominated by the former. The strategy of revealing the full extension is not an exception to this. We show this fact by introducing the following example.
@@
Therefore, $a_{1}$ (strictly) prefers $g\left(f^{U}\left(d_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)$. On the other hand, there is no pair $\left(p_{a_{2}},p_{a_{3}}\right)$ of preference orders of $a_{2}$ and $a_{3}$ such that $a_{1}$ (strictly) prefers $g\left(f^{U}\left(e_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$. This is because, by $q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|$, for any $\left(p_{a_{2}},p_{a_{3}}\right)\in\mathcal{P}^{2}$ and any $\bar{o}\in\mathcal{O}$,
@@
We explain why $d_{a_{1}}$ strategically dominates $e_{a_{1}}$. In this example, $q_{o_{1}}+q_{o_{3}}=2<3=q_{o_{1}}+q_{o_{2}}$. That is, the copies of $o_{1}$ and $o_{2}$ are sufficient to allocate all agents, but those of $o_{1}$ and $o_{3}$ are insufficient. Thus, revealing $o_{2}$ as a second-best type does not increase the probability that $a_{1}$ is assigned to $o_{1}$. On the other hand, revealing $o_{3}$ as a second-best type increases it, when $o_{3}$ is the first-best type of another agent. This is because, to minimize the total rank, $o_{3}$ must be assigned to an agent who ranks it as their first-best type. However, since $a_{1}$ ranks $o_{2}$ and $\varnothing$ lower, minimizing the total rank requires increasing the probability that $o_{1}$ is allocated to $a_{1}$. Thus, revealing the full extension may be strategically dominated by another ODS when an unacceptable type (in this case $o_{3}$) is less preferable than the other (in this case $o_{2}$) but the number of copies of the former is less than that of the latter.
Note that this fact does not contradict a result of Featherstone (2020, Proposition 13). In their model, they assume that the agents have no information on the capacities of object types. Revealing the full extension can be strategically dominated by another ODS only when the capacities of the unacceptable types are different and agents know them. Thus, the strategic domination in Example 3 does not occur in the low-information situation discussed by Featherstone (2020).
Next, we show that if some agent adopts an ODS, then the result may be wasteful and thus inefficient.
@@
Suppose that there is $a\in\mathcal{A}$ whose true preference satisfies (3) for some $o,o^{\prime}\in\mathcal{O}$. Then, there are some ODS of revealing $d_{a}$ and some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $R\left(\varnothing,p_{a^{\prime}}\right)=\left|\mathcal{O}\right|$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$ and $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ is wasteful.
By Theorem 2, for an agent $a$ with $r_{a}$ that satisfy (3) for some $o,o^{\prime}\in\mathcal{O}$, an ODS dominates truth-telling. Moreover, by Proposition 3, in the same case, the strategy may make the assignment wasteful. In an ODS, an agent reveals actually unacceptable types as acceptable types. Then, the agent may be assigned to an actually unacceptable type and refuse it. Since the type may be desirable for another agent, in that case, the assignment is wasteful.
By Theorems 1 and 2 and Proposition 3, if the uniform rank-minimizing mechanism is employed, the agents may have a strong incentive to manipulate their preferences, and these manipulations may result in an inefficient assignment. Thus, next, we consider another fair rank-minimizing mechanism such that, contrary to $f^{U}$, if it is employed as the mechanism in the second stage, then for any agent, truth-telling is not strategically dominated by any other strategy.
We let $f^{M}$ be the modified uniform rank-minimizing mechanism such that if $p$ satisfies
@@
To understand this mechanism, we reconsider Example 1. First, suppose $q_{o_{1}}=q_{o_{2}}=1$. If $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$, $\left(\alpha,\alpha^{\prime},\alpha\right)$ or $\left(\alpha,\alpha,\alpha^{\prime}\right)$, then $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ satisfies (i), (ii) and (iii). In these cases, the student who chooses $\alpha^{\prime}$ is corresponding to $a,$ another student is corresponding to $a^{\prime}$, and $o_{1}$ is corresponding to $o$. The rank-minimizing mechanism $f^{\ast}$ introduced in Example 1 is the modified uniform rank-minimizing mechanism, because a student who chooses $\alpha^{\prime}$ obtains $o_{1}$ with probability 00 and another student obtains $o_{1}$ with probability $1$
On the other hand, suppose $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$. Even if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$, $\left(\alpha,\alpha^{\prime},\alpha\right)$ or $\left(\alpha,\alpha,\alpha^{\prime}\right)$, this does not satisfy (ii). This is because in this case, $p_{a^{\prime}}=\alpha$ and $l=3,$ but
$\sum\limits_{\bar{o}:R\left(\bar{o},p_{a^{\prime}}\right)<R\left(\varnothing,p_{a^{\prime}}\right)}q_{\bar{o}}=q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|.$
Therefore, (ii) is not satisfied with $p_{a^{\prime}}=\alpha$ in the case of $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$.
@@
The modified uniform rank-minimizing mechanism $f^{M}$ satisfies ETE.
The strategic issues, mentioned in Theorems 1 and 2, do not arise under this mechanism $f^{M}$.
First, we explain why the strategic issue mentioned in Example 2 does not arise in the modified uniform rank-minimizing mechanism. If $r_{a}=p_{a^{\prime}}$ where $p_{a^{\prime}}$ is defined in (ii), then $\left(d_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) where $d_{a}\in\mathcal{D}_{a}$. This implies that for some $p_{-a}$, if agent $a$ adopts an ODS instead of truth-telling and $f^{M}$ is employed, then $a$ is worse off.
To see this fact, we revisit Example 2. In this example, $\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\beta,\beta^{\prime\prime},\beta^{\prime\prime}\right)$ satisfies (i), (ii) and (iii), but $\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\beta^{\prime\prime},\beta^{\prime\prime},\beta^{\prime\prime}\right)$ does not. Thus,

### v6

@@ 3 Rank-minimizing Mechanism
That is, if a rank-minimizing mechanism is employed and $a_{1}$ chooses either $\alpha$ or $\alpha^{\prime}$, then $a_{1}$ must have either $o_{1}$ or $o_{2}$. Thus, the only strategic interest of $a_{1}$ is increasing the probability to obtain $o_{1}$. If $f^{\ast}$ is employed, then $\alpha^{\prime}=r_{a_{1}}$ is dominated by $\alpha$ for $a_{1}$ due to (2). Therefore, we conclude that even if a rank-minimizing mechanism that satisfies weak ETE is employed, truth-telling may be dominated by some other strategy. Note that in the same case, $f^{\ast}$ does not satisfy ETE. First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing,$ as long as the assignment is not wasteful. Thus, $\alpha$ and $\alpha^{\prime}$ are essentially the same in this case. However, $\left(f^{\ast}\left(\alpha^{\prime},\alpha,\alpha\right)\right)_{a_{1}o_{1}}=0,$ but $\left(f^{\ast}\left(\alpha^{\prime},\alpha,\alpha\right)\right)_{a_{1}o_{2}}>0$ or $\left(f^{\ast}\left(\alpha^{\prime},\alpha,\alpha\right)\right)_{a_{1}o_{3}}>0$ must be satisfied. Hence $f^{\ast}$ does not satisfy ETE.
@@ 4 Refusal option
In this section, we consider the following two-stage game, which is a modification of that in Section 3. In the first stage, all agents simultaneously reveal a preference order in $\mathcal{P}$ and an assignment is determined via a mechanism based on $p$. In the second stage, all agents decide whether to accept the assignment determined in the first stage. If an agent refuses their assignment, then the agent instead obtains the null object type $\varnothing$ (the outside option). That is, in this game, the null object type has two characteristics: first, there are sufficiently many copies of this type to allocate to all individuals, and second, agents can always obtain it at no cost as long as they abandon their assignment.
This two-stage game is also considered by Featherstone (2020) but is not explicitly addressed by almost other previous studies related to ours. [^4] As exceptions, Feigenbaum et al. (2020) and Afacan (2022) consider an assignment problem where, after the initial allocation is decided, if someone refuses the assigned object, it is reallocated. Moreover, Doğan and Yenmez (2019) also explicitly consider this stage. However, this stage is implicitly considered by many of them; that is, they consider an ex-post individually rationality (or more strictly an ex-post efficiency) as a requirement and if an assignment is ex-post individually rational, no agent refuses their assignment. See, for example, Bogomolnaia and Moulin (2001) on the property.
@@
$g\left(x,r\right)_{ao}$ $=$ $0\text{ for all }R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$
@@
We reconsider strategic problems of a fair rank-minimizing mechanism in this three-stage game. For a mechanism $f$ and an agent $a$, (revealing) $p_{a}^{\prime}$ is said to be weakly dominated by (revealing) $p_{a}$ with the refusal option if $a$ weakly prefers $g\left(f\left(p_{a},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$. Moreover, for $f$ and $a$, (revealing) $p_{a}^{\prime}$ is said to be (strictly) dominated by (revealing) $p_{a}$ with the refusal option if $p_{a}^{\prime}$ is weakly dominated by $p_{a}$ with the refusal option and there exists some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $g\left(f\left(p_{a},p_{-a}^{\prime}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}^{\prime}\right),r\right)$. For a mechanism $f$ and an agent $a$, (revealing) $p_{a}\in\mathcal{P}$ is said to be a (weakly) dominated strategy with the refusal option if there is some $p_{a}^{\prime}\in\mathcal{P}$ that (weakly) dominates $p_{a}$ with the refusal option.
Moreover, we say that $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ is a Nash equilibrium with the refusal option if there is no $a$ who strictly prefers $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a},p_{-a}\right),r\right)$ for some $p_{a}^{\prime}\in\mathcal{P}$.
@@
We show that, under the URM, truth-telling is dominated by another strategy when the agents have the refusal option. We define the following specific strategies. Let $\mathcal{D}_{a}\subseteq\mathcal{P}$ be a set of preferences such that $d_{a}\in\mathcal{D}_{a}$ if (i) $R\left(o,r_{a}\right)=R\left(o,d_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and (ii) $R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$. First, the ranks of the truly acceptable object types are unchanged. Second, the null object type $\varnothing$ is the least preferable one under $d_{a}\in\mathcal{D}_{a}$. We call the strategy of agent $a$ of revealing a preference in $\mathcal{D}_{a}$ an outside option demotion strategy (hereafter ODS) of agent $a$. Under an ODS, every object type that is ranked below $\varnothing$ in the agent’s true preference is instead ranked above $\varnothing$ in $d_{a}$. Note that an ODS (revealing $d_{a}\in\mathcal{D}_{a}$) increases the number of acceptable object types revealed by $a$; that is, $R\left(\varnothing,r_{a}\right)\leq R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$ is satisfied.
If $R\left(\varnothing,r_{a}\right)=\left|\mathcal{O}\right|,$ then truth-telling is the unique ODS for agent $a$. Therefore, an ODS constitutes a strategic manipulation if and only if $R\left(\varnothing,r_{a}\right)<\left|\mathcal{O}\right|;$ that is, the null object is not the worst object type for $a$.
@@
Featherstone (2020) defines the full extension of $r_{a}\$denoted by $e_{a}$ in this study such that (i) $R\left(o,r_{a}\right)=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (ii) $R\left(\varnothing,e_{a}\right)=\left|\mathcal{O}\right|$, and $R\left(o,r_{a}\right)-1=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$; that is, $e_{a}\in\mathcal{D}_{a}$ is such that all non-null object types are placed in their true preference orders. Revealing the full extension of $r_{a}$ is one of the ODSs of $a$.
@@
Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their first-best object $o_{1}$. First, under $\left(r_{a_{1}},\beta,\beta^{\prime}\right),$ $\varnothing$ is ranked lowest by $a_{2}$ and $a_{3}$ but $a_{1}$ ranks $\varnothing$ as their second-best. Thus, in any rank-minimizing mechanism, $a_{1}$ is assigned $\varnothing$ with probability 1. On the other hand, since $a_{1}$ places $\varnothing$ at the bottom of their ranking. As a result. under $\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$, $a_{1}$ is assigned $o_{1}$—the object that is truly their best—with probability $1/2$, and $\varnothing$ with probability $1/2$. Thus, in this case, $a_{1}$ is better off by choosing $d_{a_{1}}$ instead of $r_{a_{1}}$.
@@
$f^{U}\left(r_{a_{1}},\beta,\beta\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1/2&1/2&0\\ 1/2&1/2&0\end{array}\right),$ (6)
$f^{U}\left(d_{a_{1}},\beta,\beta\right)=\left(\begin{array}[]{ccc}1/3&1/3&1/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$ (10)
@@
$g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$ (11)
@@
$\text{and }g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)_{a^{\prime}\varnothing}$ $=$ $\frac{1}{3}>0,$
@@
If $a_{1}$ is assigned to $o_{2}$, then either $a_{2}$ or $a_{3}$ will be assigned to $\varnothing$. However, in this case, $a_{1}$ will refuse $o_{2}$ and thus $o_{2}$ will remain unallocated. In this case, rather than leaving $o_{2}$ unallocated, it is more efficient to assign $o_{2}$ to the agent ($a_{2}$ or $a_{3}$) who would otherwise receive $\varnothing$ and instead assign $a_{1}$ to $\varnothing$ from the beginning. This adjustment ensures that $o_{2}$ is allocated efficiently. Hence revealing $d_{a_{1}}$ and refusing an unacceptable assignment may benefit $a_{1}$ but lead to a wasteful allocation, as some agents may prefer the object type that $a_{1}$ refuses.
@@
For $f^{U}$ and any agent $a$, truth-telling is weakly dominated strategy; that is, it is weakly dominated by any ODS with the refusal option.
@@
Theorems 1 and 2 do not imply that an ODS is a dominant strategy. However, if there is $a\in\mathcal{A}$ whose true preference satisfies (12) for some $o,o^{\prime}\in\mathcal{O}$, since truth-telling is a dominated strategy, $a$ has a strong incentive to manipulate its preference ranking.
As stated earlier, multiple ODSs may exist. Even if an ODS dominates truth-telling, another ODS that does not dominate truth-telling may also exist. In such a case, the latter ODS is dominated by the former. The strategy of revealing the full extension is not an exception to this. We show this fact by introducing the following example.
@@
Therefore, $a_{1}$ (strictly) prefers $g\left(f^{U}\left(d_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)$ and $r_{a_{1}}$ is dominated by $d_{a_{1}}$.
On the other hand, there is no pair $\left(p_{a_{2}},p_{a_{3}}\right)$ of preference orders of $a_{2}$ and $a_{3}$ such that $a_{1}$ (strictly) prefers $g\left(f^{U}\left(e_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$. This is because, by $q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|$, for any $\left(p_{a_{2}},p_{a_{3}}\right)\in\mathcal{P}^{2}$ and any $\bar{o}\in\mathcal{O}$,
@@
Thus, $e_{a_{1}}$ is also dominated by $d_{a_{1}}$.
We intuitively explain why $d_{a_{1}}$ dominates $e_{a_{1}}$. In this example, $q_{o_{1}}+q_{o_{3}}=2<3=q_{o_{1}}+q_{o_{2}}$. That is, the copies of $o_{1}$ and $o_{2}$ are sufficient to allocate all agents, but those of $o_{1}$ and $o_{3}$ are insufficient. Thus, revealing $o_{2}$ as a second-best type does not increase the probability that $a_{1}$ is assigned to $o_{1}$. On the other hand, revealing $o_{3}$ as a second-best type increases it, when $o_{3}$ is the first-best type of another agent. This is because, to minimize the total rank, $o_{3}$ must be assigned to an agent who ranks it as their first-best type. However, since $a_{1}$ ranks $o_{2}$ and $\varnothing$ lower, minimizing the total rank requires increasing the probability that $o_{1}$ is allocated to $a_{1}$. Thus, revealing the full extension may be dominated by another ODS when an unacceptable type (in this case $o_{3}$) is less preferable than the other (in this case $o_{2}$) but the number of copies of the former is less than that of the latter.
Note that this fact does not contradict a result of Featherstone (2020, Proposition 13). In their model, they assume that the agents have no information on the capacities of object types. Revealing the full extension can be dominated by another ODS only when the capacities of the unacceptable types are different and agents know them. Hence the strategic domination in Example 3 does not occur in the low-information situation discussed by Featherstone (2020).
Next, we show that if an agent adopts an ODS, the resulting assignment may be wasteful and thus inefficient. Moreover, such an outcome can arise in a Nash equilibrium.
@@
There exist $r\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and $d\in\mathcal{P}^{\left|\mathcal{A}\right|}$ such that $d,$ the profile in which all agents adopt ODSs, is a Nash equilibrium with the refusal option under $f^{U}$. Moreover, if there are $o,o^{\prime}\in\mathcal{O}$ such that $q_{o}+q_{o^{\prime}}<\left|\mathcal{A}\right|$, then such profiles $r\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and $d\in\mathcal{P}^{\left|\mathcal{A}\right|}$ can also yield a wasteful assignment; that is, $g\left(f^{U}\left(d\right),r\right)$ is wasteful.
We briefly show this by revisiting Example 2. Let
$r=\left(r_{a_{1}},r_{a_{2}},r_{a_{3}}\right)=\left(\beta^{\prime\prime},\beta,\beta\right)\text{ and }d=\left(d_{a_{1}},d_{a_{2}},d_{a_{3}}\right)=\left(\beta,\beta,\beta\right).$
Then, $g\left(f^{U}\left(d\right),r\right)$ is given by (11), which is a wasteful assignment. Moreover, $d=\left(\beta,\beta,\beta\right)$ is a Nash equilibrium; that is, for each agent $a$, $g\left(f^{U}\left(d\right),r\right)$ is not first-order stochastically dominated by $g\left(f^{U}\left(p_{a},d_{-a}\right),r\right)$ for any $p_{a}\in\mathcal{P}$. [^5] Since we consider the first-order stochastic dominance relation, our notion of Nash equilibrium is weaker than the standard one. Nevertheless, suppose that both $a_{2}$ and $a_{3}$ prefer the assignment such that each object $o\in\left\{o_{1},o_{2},\varnothing\right\}$ assigns with the probability $1/3$ to the assignment such that $o_{2}$ assigns with probability $1$. This assumption ensures that neither $a_{2}$ nor $a_{3}$ has an incentive to change $\beta$ and $\beta$ to $(o_{2},o_{1},\varnothing)$ or $(o_{2},\varnothing,o_{1})$. Under this condition, $d=\left(\beta,\beta,\beta\right)$ is a Nash equilibrium even in the standard sense.
By Proposition 3, if agents adopt an ODS, then the result may be wasteful. Moreover, such a strategy profile can be a Nash equilibrium. Therefore, under the URM, a profile of ODS are stably adopted in place of truth-telling and that such strategies may result in inefficient outcomes. This suggests a significant drawback of the URM.
Therefore, we next consider an alternative fair rank-minimizing mechanism which, unlike $f^{U}$, guarantees that for any agent, truth-telling is not dominated by any other strategy when it is used as the mechanism in the second stage.
We let $f^{M}$ be the modified uniform rank-minimizing mechanism (hereafter modified URM) such that if $p$ satisfies
@@
Under this mechanism, the outcome is equivalent to that of the URM, except when a specific preference profile satisfying conditions (i)–(iii) is realized. The specific profile characterized by conditions (i)–(iii) is one in which: (i) a particular agent denoted by $a$ ranks object $o$ first and the outside option at position three or lower; (ii) there are $q_{o}$ other agents whose preferences match $a$’s for the top l1l-1 positions, rank the outside option exactly $l$th, and for whom the total capacity of objects ranked above $l$ is insufficient for all agents; and (iii) if other agents remain, all of them rank the outside option first.
@@
The modified URM satisfies ETE.
Then, we explain why the strategic issues, mentioned in Theorems 1 and 2, do not arise under this mechanism $f^{M}$. To ensure that truth-telling is not dominated by any ODS, the mechanism must be such that, for every agent and ODS, there exists a strategy profile of the other agents under which the agent is not assigned to any of the more preferred object types. Suppose that the true preference of $a$ satisfies $r_{a}^{1}=o$ and $r_{a}^{2}=\varnothing$. Given $p_{-a}$ satisfying conditions (ii) and (iii), condition (i) is not satisfied if agent $a$ chooses truth-telling, but it is satisfied if $a$ adopts an ODS. Therefore, when agent $a$ adopts an ODS, the probability of being assigned their most preferred object $o$ becomes zero, whereas under truth-telling, this probability is positive.
To understand this point more, we revisit Example  2. In this example, $\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\beta,\beta^{\prime\prime},\beta^{\prime\prime}\right)$ satisfies (i), (ii) and (iii), but $\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\beta^{\prime\prime},\beta^{\prime\prime},\beta^{\prime\prime}\right)$ does not. Thus,